Your credit card has a current balance of $4,965.20. This balance is accruing interest at a nominal rate of 24.0% compounded monthly. Suppose you plan on spending an additional $1,000 at the end of next year (end of month 12) and $1,000 at the end of the following year (end of month 24) on this same card. What uniform end-of-month payments over the next 36 months would be required to pay off the existing debt along with the additional new purchases?
Answer #1

Answer :

Given interest rate =24% per annum

So monthly Interst rate =24%/12=2%

Pv of credit card spending =4965.20+1000/(1+2%)^12+1000/(1+2%)^24

Pv of credit card spending =$6906.76

Let x be the monthly payment for credit card

Pv of this monthly payment =x*(1-(1+r)^-n)/r

r=2%

n=36 months

Pv of monthly payment =x*(1-(1+2%)^-36)/2%=25.48x

So pv of monthly payment = pv of credit card spending

25.48x=6906.76

X=$271